OpenStudy (anonymous):

How many solutions does the equation 5x + 10 = 5x − 8 have? Zero One Two Infinitely many

OpenStudy (anonymous):

@Baby_Bear69

Nnesha (nnesha):

solve for x if you get equal sides like 5=5 thats mean Infinitely many solutions and if you get x= something (then you will have one solution) and both sides are not equal like 4=9 then there is no solution

OpenStudy (anonymous):

one?

Nnesha (nnesha):

how ?

OpenStudy (anonymous):

Zero

OpenStudy (baby_bear69):

It would be zero because theres no way 5x+10 could equal 5x - 8

OpenStudy (anonymous):

one more?

OpenStudy (anonymous):

5x+10=5x−8 Subtract 5x from both sides. 5x+10−5x=5x−8−5x 10=−8 Subtract 10 from both sides. 10−10=−8−10 0=−18 NO SOLUTION! :D

OpenStudy (anonymous):

wait what..

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @bby.nessa i was right yaayyy thankssss \(\color{blue}{\text{End of Quote}}\) no

OpenStudy (anonymous):

oh..

OpenStudy (baby_bear69):

The answer is zero and you had said 1 @bby.nessa

OpenStudy (ali2x2):

lol nnesha you killed the moment

Nnesha (nnesha):

and all of you who just gave out the direct answer should read this http://openstudy.com/code-of-conduct and welcome to openstudy

OpenStudy (anonymous):

haha fr

OpenStudy (anonymous):

.......

OpenStudy (baby_bear69):

I didn't give a direct answer, i explained .-.

OpenStudy (anonymous):

no chill..

OpenStudy (anonymous):

What is the value of x in the equation 2(2x − 5 − 4) = 160 − 17? 40.25 62.25 80.25 143.0

Nnesha (nnesha):

let her think about it for a sec :)

OpenStudy (anonymous):

2(2x−5−4)=160−17 (2)(2x)+(2)(−5)+(2)(−4)=160+−17(Distribute) 4x+−10+−8=160+−17 (4x)+(−10+−8)=(160+−17)(Combine Like Terms) 4x+−18=143 4x−18=143 Add 18 to both sides. 4x−18+18=143+18 4x=161 Divide both sides by 4. 4x/4=16/14 x= ? U can finish it up.

Nnesha (nnesha):

still...

OpenStudy (baby_bear69):

@Excellency that's what i was trying to write but you saved me. thanks

OpenStudy (anonymous):

Np.

OpenStudy (anonymous):

ok i got it thanks